# Finite–slope universal Rankin–Selberg $p$-adic $L$-function

### v1

@@ Abstract
In this article we construct a finite–slope universal Rankin–Selberg $p$-adic $L$-function attached to a half–ordinary universal deformation family of $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ Galois representations. More precisely, starting from residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ of tame level $1$ satisfying Hypothesis 3.1 of Loeffler [13], we consider the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of [13], where the first factor varies in the ordinary Hida deformation and the second factor in the unrestricted universal deformation space. Fixing a classical point $x_{0}$ on a suitable parabolic eigenvariety for which the first factor is non–ordinary but of small finite slope, we use Liu’s global triangulation theorem, Kedlaya–Pottharst–Xiao’s $(\varphi,\Gamma)$–cohomology, and the Perrin–Riou–Loeffler–Zerbes regulator formalism to attach to the resulting Panchishkin family $(V_{U},V_{U}^{+})$ over a neighbourhood $U$ of $x_{0}$ a “big logarithm”
$\mathcal{L}_{V_{U},V_{U}^{+}}\colon H^{1}_{\mathrm{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)\cong\mathcal{O}(U\times\mathscr{W}),$
interpolating the Bloch–Kato dual exponentials at all classical points in $U$.
Combining this family regulator with the universal Beilinson–Flach Euler system of Loeffler-Zerbes[21], as interpreted in the universal deformation setting in [15], we define a rigid–analytic function
@@
and prove that $L_{p}^{\mathrm{fs}}$ satisfies the expected interpolation formula for all Deligne–critical values of the complex Rankin–Selberg $L$-functions $L(f_{x}\otimes g_{x},s)$ at classical points $(x,\kappa)\in U\times\mathscr{W}$. In particular, under mild crystalline and non–criticality hypotheses at $p$, we verify the finite–slope analogue of Loeffler’s Conjecture 2.8 in the concrete $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ half–ordinary universal deformation setting of [13].
The main new features of this work are: (1) the construction of a Perrin–Riou regulator for a genuinely finite–slope Panchishkin family over a parabolic eigenvariety in the universal deformation context, and (2) the resulting finite–slope universal Rankin–Selberg $p$-adic $L$-function, which simultaneously extends the ordinary universal results of Loeffler and Hao–Loeffler [13, 15] and the finite–slope eigencurve constructions of Loeffler’s earlier three–variable Rankin–Selberg $p$-adic $L$-functions. This provides the first verification of the universal finite–slope conjectural picture in the Rankin–Selberg case and offers a conceptual bridge between ordinary Hida-theoretic constructions and finite–slope Coleman–style $p$-adic $L$-functions.
@@ 1 Introduction
1. Global goal and unsolved question
We fix a prime $p\geq 3$ and an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. Let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and $G_{\mathbb{Q},\{p\}}\subset G_{\mathbb{Q}}$ denote the Galois group of the maximal extension of $\mathbb{Q}$ unramified outside $p$.
@@ 1. Global goal and unsolved question
be continuous, odd, absolutely irreducible Galois representations arising from cuspidal newforms $f_{0},g_{0}$ of weights $k_{1},k_{2}\geq 2$ and tame level $N=1$ (unramified outside $p$), so that Hypothesis 3.1 of [13] applies. In particular (see [13, Hyp. 3.1]), we assume that:
  • both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ are modular of level $1$ and weight $\geq 2$;
@@
We henceforth keep this tame level $N=1$ fixed throughout.
@@
    be the corresponding ordinary eigenvariety, as in [13, §3.2, Prop. 3.14], building on the ordinary families of Hida and the $R=\mathbb{T}$ theorems of Mazur, Böckle and Emerton.
  • Let $R_{2}$ be the universal deformation ring of $\bar{\rho}_{2}$ (for deformations unramified outside $p$) and let $(R_{2},V_{2})$ be the Galois deformation family considered in [13, Def. 3.3, Thm. 3.4] and in [14, 15]. We denote
@@
where $V_{1}^{\mathrm{ord}}$ is the universal ordinary deformation of $\bar{\rho}_{1}$ and $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ is the rank–$1$ ordinary submodule at $p$. More precisely, there is a rank–$r$ locally free RR–submodule
@@
Loeffler’s Conjecture 2.8 [13, Conj. 2.8] predicts the existence of a rank 0 Euler system (i.e. a $p$-adic $L$-function) attached to $(R,V,V^{+})$; in the ordinary setting this is now known by [15, Thm. 3.5], building on Urban’s construction via nearly overconvergent forms [7] and Hida’s earlier work on Rankin–Selberg pp–adic LL–functions [1, 2].
In [13, §5.5] Loeffler sketches a finite-slope analogue of this conjecture on the Coleman–Mazur eigencurve, and in [13, §6] he formulates related conjectures over big parabolic eigenvarieties (see also [9, 12]). However, beyond very special cases for $\operatorname{GL}_{2}$–families and the rank-one case of [11], no general finite-slope result is currently known.
Unsolved question (finite-slope universal Rankin–Selberg). Let $\mathscr{E}_{1}$ be the Coleman–Mazur eigencurve (or more generally the big parabolic eigenvariety of Barrera Salazar–Williams) attached to $\bar{\rho}_{1}$ [23, 9], and consider the product eigenvariety
@@
together with the natural weight map to the pp–adic weight space $\mathscr{W}$ and the global Galois family $(V,V^{+})$ coming from [13, Ex. 3.17].
Given a classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ of cuspidal eigenforms with $f_{0}$ of non-ordinary finite slope at $p$, does there exist a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on a neighbourhood $U$ of $x_{0}$ whose specialisations interpolate all Deligne-critical Rankin–Selberg values $L(f_{x}\otimes g_{x},s_{x})$ for classical points $x\in U$?
This is a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ instance of the finite-slope variants of [13, Conj. 2.8] over parabolic eigenvarieties.
To treat this question we need a precise description of the eigenvariety $\mathscr{E}$ near $x_{0}$. We recall the relevant geometric input from [9, Han17], in a form adapted to our setting.
Lemma 1 (Good neighbourhood on the parabolic eigenvariety).
Let $\mathscr{E}$ be the parabolic eigenvariety of [9] with weight space $\mathscr{W}$ and weight map $w:\mathscr{E}\to\mathscr{W}$. Let $x_{0}\in\mathscr{E}$ be a point corresponding to a cuspidal automorphic representation of regular (cohomological) weight, and fix a pp–refinement which is QQ–non-critical in the sense of [9, Def. 5.13]. Assume that the derived group $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series.
Then there exists an affinoid neighbourhood $U\subset\mathscr{E}$ of $x_{0}$ such that:
@@
    the restriction $w|_{U}:U\to w(U)\subset\mathscr{W}$ is finite étale; and
@@
On the other hand, $\mathscr{E}$ is constructed as a closed subspace of the universal eigenvariety of Hansen [Han17]. By [Han17, Thm. 4.5.1(i)], applied to the corresponding classical, non-critical interior point of the universal eigenvariety, every irreducible component of the latter has dimension equal to the weight space. Since $\mathscr{E}$ is a closed subspace of this universal eigenvariety, every irreducible component $V$ of $\mathscr{E}$ through $x_{0}$ has dimension at most $\dim\mathscr{W}$. Thus $\dim V=\dim\mathscr{W}$ for each component $V$ passing through $x_{0}$.
@@
The weight map $w:\mathscr{E}\to\mathscr{W}$ is finite by [9, Thm. 5.4]. Since the source and the target have the same local dimension at $x_{0}$, the morphism $w$ is unramified at $x_{0}$. For finite morphisms between equidimensional rigid analytic spaces, finiteness and unramifiedness imply that $w$ is étale at $x_{0}$ and that $\mathscr{E}$ is smooth at $x_{0}$. The smooth locus of $\mathscr{E}$ and the étale locus of $w$ are open subsets of $\mathscr{E}$, so we may choose an affinoid neighbourhood $U$ of $x_{0}$ contained in their intersection. By the density of classical points on each component through $x_{0}$, the intersection of $U$ with the classical cuspidal locus is Zariski dense in $U$. This proves the lemma. ∎
Remark 1 (On the discrete-series hypothesis in Lemma 1).
@@
which is well known to admit discrete series. Thus the hypothesis that $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series in Lemma 1 is automatically satisfied in our setting.
The only role of this hypothesis is to allow us to invoke the eigenvariety results of [9, Han17], in particular the dimension and density statements of [9, Prop. 5.12, Cor. 5.16] and [Han17, Thm. 4.5.1(i)], in order to construct an affinoid neighbourhood $U$ of $x_{0}$ which is smooth, finite étale over weight space, and has Zariski-dense classical cuspidal locus. All later arguments only use the existence of such a neighbourhood $U$ and do not appeal directly to the discrete-series condition.
2. Target main theorem (precise special case)
Fix a finite-slope classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ with:
  • $f_{0}$ a $p$-stabilised newform of weight $k_{1}\geq 2$ and level prime to $p$, with $U_{p}$–eigenvalue $\alpha_{f_{0}}$ of slope
@@ 2. Target main theorem (precise special case)
  • $g_{0}$ the specialisation at a classical arithmetic point of the universal deformation family $(R_{2},V_{2})$, of weight $k_{2}\geq 2$, such that $g_{0}$ is ordinary at $p$ (or more generally satisfies a suitable Panchishkin condition as in [13, Def. 2.1, Def. 2.3] and [15, §2]).
Let $U\subset\mathscr{E}$ be the affinoid neighbourhood of $x_{0}$ given by Lemma 1. Thus $\mathscr{E}$ is smooth at every point of $U$, the weight map $w$ is finite étale on $U$, and classical cuspidal points are Zariski dense in $U$.
@@
  • the eigenvariety is smooth on $U$ and the weight map $w\colon U\to w(U)\subset\mathscr{W}$ is finite étale (Lemma 1);
@@
  • the Panchishkin submodule $V^{+}$ extends as a rank–$r$ subbundle of the relative $(\varphi,\Gamma)$–module attached to $V$ over $U$, as in [13, Def. 2.1, Def. 2.3, Ex. 3.17] and Liu’s global triangulation theorem [20, Thm. 1.8].
Conjecture 2 (Finite-slope universal Rankin–Selberg $p$-adic $L$-function).
@@
$L_{p}^{\mathrm{fs}}\;\in\;\mathscr{O}(U\times\mathscr{W})$
on $U$ times the $p$-adic cyclotomic weight space $\mathscr{W}$ such that for each classical point $(x,\kappa)\in U\times\mathscr{W}$ corresponding to a pair $(f_{x},g_{x})$ and an integer critical value $s=\kappa(x)$ of $L(f_{x}\otimes g_{x},s)$ (in Deligne’s sense), we have an interpolation formula of the form
@@
Theorem 3 (Target main result of the paper).
@@
    The residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ satisfy Hypothesis 3.1 of [13] (Taylor–Wiles conditions, local conditions at $p$, and tame level $1$). In particular, both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ arise from cuspidal newforms of level $1$ and weights $\geq 2$, and we fix once and for all this tame level $N=1$ throughout. Moreover, $\bar{\rho}_{1}$ admits an ordinary refinement at $p$ in the sense of [13, Def. 3.11].
@@
    the point $x_{0}$ is crystalline at $p$ and satisfies the small-slope and non-criticality hypotheses needed for overconvergent classicality (in the sense of Coleman [22, Thm. 6.1]);
@@
    the Panchishkin condition holds for $(V,V^{+})$ on $U$ in the sense of [13, Def. 2.1, Def. 2.3] (in particular, the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of loc. cit. uniformly over $U$).
@@
Then Conjecture 2 holds for the neighbourhood $U$. In particular, there exists a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on $U\times\mathscr{W}$ satisfying the above interpolation formula for all classical points in $U$.
@@
The rest of the paper is devoted to the construction of $L_{p}^{\mathrm{fs}}$ and the proof of Theorem 3. Each step is explicitly referenced to the corresponding statements in [13, 19, 20, 16, 21, 25, 15, 14] so that the argument can be checked line by line.
3. New features of this work, obstacles, and future directions.
We conclude the introduction by explaining how Theorem 3 fits into the existing conjectural picture of universal Rankin–Selberg $p$-adic $L$-functions, why the finite–slope case treated here has remained out of reach so far, and which directions seem most promising for further developments.
3.1. Relation to Loeffler’s conjectural framework and to Hao–Loeffler.

### v2

# Towards a finite–slope universal Rankin–Selberg $p$-adic $L$-function

@@ Abstract
This article studies the finite–slope analogue of Loeffler’s conjectural framework for Rankin–Selberg $p$-adic $L$-functions in universal deformation families. Starting from residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ of tame level $1$ satisfying Hypothesis 3.1 of [13], we consider the half–ordinary Panchishkin family $(R,V,V^{+})$ of Example 3.17 of loc. cit., where the first factor varies in the ordinary Hida deformation and the second factor in the unrestricted universal deformation space.
We fix a classical point $x_{0}$ on a suitable parabolic eigenvariety for which the first factor is non–ordinary but of small finite slope. Using Liu’s global triangulation theorem, the cohomology of families of $(\varphi,\Gamma)$-modules due to Kedlaya–Pottharst–Xiao, and the Perrin–Riou / Loeffler–Zerbes regulator formalism, we attach to the resulting finite–slope Panchishkin data over a neighbourhood $U$ of $x_{0}$ a family “big logarithm”
$\mathcal{L}_{V_{U},V_{U}^{+}}\colon H^{1}_{\operatorname{Iw}}(\mathbb{Q}_{p},V_{U}^{\ast}(1))\longrightarrow\mathcal{H}(\Gamma)\widehat{\otimes}_{\mathbb{Q}_{p}}\mathcal{O}(U)\cong\mathcal{O}(U\times\mathscr{W}),$
which interpolates the Bloch–Kato dual exponentials at all classical points of $U$.
Assuming, in addition, the existence of a global Beilinson–Flach Euler system for this half–ordinary universal deformation family (formulated precisely as Conjecture 2.15 below), we apply this family regulator to the resulting Iwasawa cohomology class and obtain a rigid–analytic function
@@
which we refer to as a finite–slope universal Rankin–Selberg $p$-adic $L$-function. Under standard global and local hypotheses, and conditional on Conjecture 2.15, we show that $L_{p}^{\mathrm{fs}}$ satisfies the expected interpolation formula for all Deligne–critical values of the complex Rankin–Selberg $L$-functions $L(f_{x}\otimes g_{x},s)$ at classical points $(x,\kappa)\in U\times\mathscr{W}$.
Thus the main unconditional output of this work is the construction of a Perrin–Riou regulator for a genuinely finite–slope Panchishkin family over a parabolic eigenvariety in the universal deformation setting. The existence of the associated universal Beilinson–Flach Euler system and the resulting $p$-adic $L$-function remains conjectural; if it holds, our construction would give a finite–slope analogue of the Rankin–Selberg case of Loeffler’s Conjecture 2.8 in [13] for a concrete $\mathrm{GL}_{2}\times\mathrm{GL}_{2}$ half–ordinary universal deformation family.
@@ 1 Introduction
1.1 Global goal and formulation of the problem
We fix a prime $p\geq 3$ and an algebraic closure $\overline{\mathbb{Q}}$ of $\mathbb{Q}$. Let $G_{\mathbb{Q}}:=\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and let $G_{\mathbb{Q},\{p\}}\subset G_{\mathbb{Q}}$ denote the Galois group of the maximal extension of $\mathbb{Q}$ unramified outside $p$.
@@ 1.1 Global goal and formulation of the problem
be continuous, odd, absolutely irreducible Galois representations arising from cuspidal newforms $f_{0},g_{0}$ of weights $k_{1},k_{2}\geq 2$ and tame level $N=1$ (unramified outside $p$), so that Hypothesis 3.1 of [13] applies. In particular, we assume that:
  1. (a)
    both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ are modular of level $1$ and weight at least $2$;
  2. (b)
@@
We keep this tame level $N=1$ fixed throughout.
@@
be the corresponding ordinary eigenvariety, as in [13, §3.2, Prop. 3.14], building on the ordinary families of Hida and the $R=\mathbb{T}$ theorems of Mazur, Böckle and Emerton. Let $R_{2}$ be the universal deformation ring of $\bar{\rho}_{2}$ (for deformations unramified outside $p$) and let $(R_{2},V_{2})$ be the Galois deformation family considered in [13, Def. 3.3, Thm. 3.4] and in [14, 15]. We denote
@@
where $V_{1}^{\mathrm{ord}}$ is the universal ordinary deformation of $\bar{\rho}_{1}$ and $V_{1}^{\mathrm{ord},+}\subset V_{1}^{\mathrm{ord}}\big|_{G_{\mathbb{Q}_{p}}}$ is the rank one ordinary submodule at $p$. More precisely, there is a rank $r$ locally free $R$-submodule
@@
which is stable under $G_{\mathbb{Q}_{p}}$ and satisfies the rr–Panchishkin condition in the sense of [13, Def. 2.1, Def. 2.3].
@@
Loeffler’s Conjecture 2.8 in [13] predicts the existence of a rank 0 Euler system (i.e. a $p$-adic $L$-function) attached to $(R,V,V^{+})$. In the ordinary setting this is now known by [15, Thm. 3.5], building on Urban’s construction via nearly overconvergent forms [7] and Hida’s earlier work on Rankin–Selberg $p$-adic $L$-functions [1, 2].
In [13, §5.5] a finite-slope analogue of this conjecture is sketched on the Coleman–Mazur eigencurve, and in [13, §6] related conjectures over big parabolic eigenvarieties are formulated; see also [9, 12]. Apart from very special cases for $\operatorname{GL}_{2}$-families and the rank one case of [11], there is at present no general finite-slope result in the universal deformation setting.
We now formulate the basic question that motivates this work.
Question (finite-slope universal Rankin–Selberg). Let $\mathscr{E}_{1}$ be the Coleman–Mazur eigencurve (or more generally the big parabolic eigenvariety of Barrera Salazar–Williams) attached to $\bar{\rho}_{1}$ [23, 9], and consider the product eigenvariety
@@
together with the natural weight map to the pp–adic weight space $\mathscr{W}$ and the global Galois family $(V,V^{+})$ coming from [13, Ex. 3.17]. Given a classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ of cuspidal eigenforms with $f_{0}$ of non–ordinary finite slope at $p$, does there exist a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on a neighbourhood $U$ of $x_{0}$ whose specialisations interpolate all Deligne–critical Rankin–Selberg values $L(f_{x}\otimes g_{x},s_{x})$ for classical points $x\in U$?
This is a concrete $\operatorname{GL}_{2}\times\operatorname{GL}_{2}$ instance of the finite-slope variants of [13, Conj. 2.8] over parabolic eigenvarieties. In this paper we give a conditional positive answer, assuming a natural Euler–system conjecture for the relevant universal deformation family.
To treat this question we need a precise description of the eigenvariety $\mathscr{E}$ near $x_{0}$. We recall the relevant geometric input from [9, 29], in a form adapted to our setting.
Lemma 1.1 (Good neighbourhood on the parabolic eigenvariety).
Let $\mathscr{E}$ be the parabolic eigenvariety of [9] with weight space $\mathscr{W}$ and weight map $w:\mathscr{E}\to\mathscr{W}$. Let $x_{0}\in\mathscr{E}$ be a point corresponding to a cuspidal automorphic representation of regular (cohomological) weight, and fix a pp–refinement which is QQ–non-critical in the sense of [9, Def. 5.13]. Assume that the derived group $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series. Then there exists an affinoid neighbourhood $U\subset\mathscr{E}$ of $x_{0}$ such that:
@@
    the restriction $w|_{U}:U\to w(U)\subset\mathscr{W}$ is finite étale;
@@
On the other hand, $\mathscr{E}$ is constructed as a closed subspace of the universal eigenvariety of Hansen [29]. By [29, Thm. 4.5.1(i)], applied to the corresponding classical, non-critical interior point of the universal eigenvariety, every irreducible component of the latter has dimension equal to the weight space. Since $\mathscr{E}$ is a closed subspace of this universal eigenvariety, every irreducible component $V$ of $\mathscr{E}$ through $x_{0}$ has dimension at most $\dim\mathscr{W}$. Thus $\dim V=\dim\mathscr{W}$ for each component $V$ passing through $x_{0}$.
@@
The weight map $w:\mathscr{E}\to\mathscr{W}$ is finite by [9, Thm. 5.4]. Since the source and the target have the same local dimension at $x_{0}$, the morphism $w$ is unramified at $x_{0}$. For finite morphisms between equidimensional rigid analytic spaces, finiteness and unramifiedness imply that $w$ is étale at $x_{0}$ and that $\mathscr{E}$ is smooth at $x_{0}$. The smooth locus of $\mathscr{E}$ and the étale locus of $w$ are open subsets of $\mathscr{E}$, so we may choose an affinoid neighbourhood $U$ of $x_{0}$ contained in their intersection. By the density of classical points on each component through $x_{0}$, the intersection of $U$ with the classical cuspidal locus is Zariski dense in $U$. ∎
Remark 1.2 (On the discrete-series hypothesis in Lemma 1.1).
@@
which is well-known to admit discrete series. Thus the hypothesis that $G_{\mathrm{der}}(\mathbb{R})$ admits discrete series in Lemma 1.1 is automatically satisfied in our setting. The only role of this hypothesis is to allow us to invoke the eigenvariety results of [9, 29], in particular the dimension and density statements of [9, Prop. 5.12, Cor. 5.16] and [29, Thm. 4.5.1(i)], in order to construct an affinoid neighbourhood $U$ of $x_{0}$ which is smooth, finite étale over weight space, and has Zariski-dense classical cuspidal locus. All later arguments only use the existence of such a neighbourhood $U$ and do not appeal directly to the discrete-series condition.
1.2 Target main result
We now specialise to the Rankin–Selberg setting of interest. Fix a finite-slope classical point $x_{0}\in\mathscr{E}$ corresponding to a pair $(f_{0},g_{0})$ with the following properties:
  1. (a)
    $f_{0}$ is a $p$-stabilised newform of weight $k_{1}\geq 2$ and level prime to $p$, with $U_{p}$–eigenvalue $\alpha_{f_{0}}$ of slope
@@ 1.2 Target main result
  2. (b)
    $g_{0}$ is the specialisation at a classical arithmetic point of the universal deformation family $(R_{2},V_{2})$, of weight $k_{2}\geq 2$, such that $g_{0}$ is ordinary at $p$, or more generally satisfies a suitable Panchishkin condition as in [13, Def. 2.1, Def. 2.3] and [15, §2].
Let $U\subset\mathscr{E}$ be the affinoid neighbourhood of $x_{0}$ given by Lemma 1.1. Thus $\mathscr{E}$ is smooth at every point of $U$, the weight map $w$ is finite étale on $U$, and classical cuspidal points are Zariski dense in $U$.
@@
  1. (i)
    the eigenvariety is smooth on $U$ and the weight map $w\colon U\to w(U)\subset\mathscr{W}$ is finite étale (Lemma 1.1);
  2. (ii)
@@
  3. (iii)
[…]
The following conjecture is the finite-slope analogue of Loeffler’s universal Rankin–Selberg $p$-adic $L$-function in this neighbourhood.
Conjecture 1.3 (Finite-slope universal Rankin–Selberg $p$-adic $L$-function).
@@
$L_{p}^{\mathrm{fs}}\in\mathcal{O}(U\times\mathscr{W})$
on $U$ times the $p$-adic cyclotomic weight space $\mathscr{W}$ such that for each classical point $(x,\kappa)\in U\times\mathscr{W}$ corresponding to a pair $(f_{x},g_{x})$ and an integer critical value $s=\kappa(x)$ of $L(f_{x}\otimes g_{x},s)$ in Deligne’s sense, one has an interpolation formula
@@
We now state the main theorem of the paper in a precise special case. It is conditional on an Euler–system conjecture which will be formulated in Conjecture 2.15.
Theorem 1.4 (Main result, conditional on Conjecture 2.15).
@@
    The residual representations $\bar{\rho}_{1},\bar{\rho}_{2}$ satisfy Hypothesis 3.1 of [13] (Taylor–Wiles conditions, local conditions at $p$, and tame level $1$). In particular, both $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ arise from cuspidal newforms of level $1$ and weights at least $2$, and we fix once and for all this tame level $N=1$ throughout. Moreover, $\bar{\rho}_{1}$ admits an ordinary refinement at $p$ in the sense of [13, Def. 3.11].
@@
    The point $x_{0}$ is crystalline at $p$ and satisfies the small-slope and non-criticality hypotheses needed for overconvergent classicality (in the sense of Coleman [22, Thm. 6.1]).
@@
    The Panchishkin condition holds for the local Rankin–Selberg representation at $p$ on $U$ in the sense of [13, Def. 2.1, Def. 2.3] (in particular, the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of loc. cit. uniformly over $U$), so that the triangulation locus of Liu contains $U$.
@@
Assume moreover the Euler–system Conjecture 2.15 for the half–ordinary universal deformation family $(R,V,V^{+})$ of [13, Ex. 3.17]. Then Conjecture 1.3 holds for the neighbourhood $U$. In particular, under (H1)–(H3) and Conjecture 2.15 there exists a finite-slope universal $p$-adic Rankin–Selberg $L$-function $L_{p}^{\mathrm{fs}}$ on $U\times\mathscr{W}$ satisfying the above interpolation formula for all classical points in $U$.
@@
The rest of the paper is devoted to the construction of the family regulator $\mathcal{L}_{V_{U},V_{U}^{+}}$ and, under Conjecture 2.15, to the construction and interpolation properties of $L_{p}^{\mathrm{fs}}$. Each step is explicitly referenced to the corresponding statements in [13, 19, 20, 16, 21, 25, 15, 14] so that the argument can be checked in detail.
1.3 Discussion and relation to previous work
In this subsection we give some context for Theorem 1.4, explain how the finite-slope case fits into existing conjectures, and briefly indicate why it is technically more delicate.
Relation to Loeffler’s conjectural framework and to Hao–Loeffler